# Embodying Control in Soft Multistable Grippers from morphofunctional co-design

### v1

@@ A.2.3 Parameter Tuning
To determine each of our lumped model’s parameters, the DPF’s fundamental unit cell is isolated into two different parts (see Fig. S3a). The bistable unit fitted as a nonlinear spring following Eq. 1 and the strain limiting layer represented by a liner spring to capture stretching and two rotational springs to capture the bending rigidity. Each of the constants for these springs needs to be tuned for different geometrical cases to produce an accurate model for inverse co-design. Given Eq. 1, the parameters to be determined for the nonlinear spring are $k_{b}$, $\alpha$, and $d$. These parameters are functions of the dome unit’s mechanical behavior (stiffness and energy barrier) as well as the interaction between the dome and the wall of the air chamber of the subsequent unit cell.
The procedure to find each constant as a function of the geometrical parameters of the dome unit cell is as follows. First, different finite element simulations of the unit cell are utilized to determine the best-fitted values for $k_{b}$, $\alpha$, and $d$ for different geometrical cases ($H$ and $t$, same as show in Fig. S1). Eq. 1 is fitted to the FE data using the nonlinear least square method to obtain the overall behavior of these parameters in terms of $H$ and $t$. Results for these fitted parameters can be observed in Fig. S3b, and the energy-displacement curve with the fitted parameters can be observed in Fig. S4a.
@@
Figure S3: a) DPF fundamental unit cell isolated into different lumped parameter spring elements. (Non-linear + linear + rotational springs) b) Non-linear spring parameter behavior over different dome heights $H$ and dome thickness $t$.
@@
Once the values are obtained, we use Automatic Relevance Determination(ARD) [48] regression to determine an expression for $k_{b}$, $\alpha$, and $d$ as functions of the dome height ($H$), dome thickness ($t$), and dome curvature ($R$). By using ARD, we can automatically determine the relevance of each feature in our model. Given the main task of inverse design of multistable actuators, a general and scalable model is needed to cover different design cases. We create the model by considering three different non-dimensional relations:
@@
where superscripts $n$ and $m$ represent the maximum degree of the polynomial use for the regression and subscripts $i$ and $j$ represent each of the non-dimensional numbers. We can fully represent this regression as:
$\begin{bmatrix}\alpha(H_{1},t_{1},R_{1})\\ \alpha_{2}(H_{2},t_{2},R_{2})\\ \vdots\\ \alpha_{k}(H_{k},t_{k},R_{k})\end{bmatrix}=\begin{bmatrix}\pi_{i}(H_{1},t_{1},R_{1})&\pi_{i}^{2}(H_{1},t_{1},R_{1})&\ldots&\pi^{n}_{i}(H_{1},t_{1},R_{1})&\pi_{i}\pi_{j}(H_{1},t_{1},R_{1})&\ldots&\pi^{n}_{i}\pi^{m}_{j}(H_{1},t_{1},R_{1})\\ \pi_{i}(H_{2},t_{2},R_{2})&\pi_{i}^{2}(H_{2},t_{2},R_{2})&\ldots&\pi^{n}_{i}(H_{2},t_{2},R_{2})&\pi_{i}\pi_{j}(H_{2},t_{2},R_{2})&\ldots&\pi^{n}_{i}\pi^{m}_{j}(H_{2},t_{2},R_{2})\\ \vdots\\ \pi_{i}(H_{k},t_{k},R_{k})&\pi_{i}^{2}(H_{k},t_{k},R_{k})&\ldots&\pi^{n}_{i}(H_{k},t_{k},R_{k})&\pi_{i}\pi_{j}(H_{k},t_{k},R_{k})&\ldots&\pi^{n}_{i}\pi^{m}_{j}(H_{k},t_{k},R_{k})\end{bmatrix}\begin{bmatrix}\xi_{1}\\ \xi_{2}\\ \vdots\\ \xi_{k}\end{bmatrix}$ (S.13)
Note that we assume that the library of terms used to create $\mathbf{C}(\pi_{1},\pi_{2},\pi_{3})$ has a sufficiently rich column space that the behavior will be represented by Eq. S.11, and it can be written as a linear combination of the weights $\xi$.
The data is split into 70% train and 30% test data sets to calculate the accuracy of our model. Observation vs prediction plot and estimation error for the three different variables can be observed in Fig. S4b (i) and (ii) respectably, where a good fit between the proposed model and the test data is observed. Moreover, we determine the precision of the distributions of the weights that allow us to determine the most relevant parameters of our regression. By establishing a threshold, the model can be further reduced by removing terms until the prediction error of the original model increases by 50% its original value. This gives a reduced model with similar prediction capabilities and the most relevant geometric relations of the problem. Our reduced model is tested against all the data (train + test), where a good agreement is found even after removing the extra terms (see Fig. S4b (iii)).
Figure S4: Parameter fitting results for $\alpha$, $k_{b}$, $d$. a) Monostable (i), metastable (ii) and bistable (iii) energy-displacement model fit vs FE simulations. b) ARD regression full model observation vs. prediction plots (test data) (i), prediction error distribution for test data, and reduced model observation vs. prediction plots (all data)
@@
$\begin{split}k_{b}=f\left(E,\pi_{1},\pi_{2},\pi_{3}\right)=&\frac{E}{{HR^{3}}}(-0.71975H^{4}+13.7556H^{3}t+H^{2}\left(0.511574R^{2}-9.43363Rt-18.4013t^{2}\right)\\ &+Ht\left(-1.55186R^{2}+39.3601Rt-46.6357t^{2}\right)+Rt^{2}(1.36789R-18.2071t))\end{split}$ (S.14)
$\begin{split}\alpha=f\left(\pi_{1},\pi_{2},\pi_{3}\right)=&\frac{0.0944903t^{3}}{H^{3}}+\frac{0.190736H^{2}}{R^{2}}+\frac{t^{2}(-0.256867R-0.290149t)}{H^{2}R}+\frac{H(-0.138165R-1.54765t)}{R^{2}}+\\ &\frac{0.679161t}{H}+\frac{t\left(0.705331R^{2}+0.707233Rt+3.87065t^{2}\right)}{R^{3}}\end{split}$ (S.15)
$d=f\left(\pi_{1},\pi_{2},\pi_{3}\right)=\frac{-1.3H^{2}Rt+2.51HR^{3}-11.34HRt^{2}+464t^{4}}{R^{3}}-H-2t-\text{U}_{\text{sep}}$ (S.16)
@@
where $H$, $t$ and $R$ are dome unit parameters, $E$ is the elastic modulus of the material and $\text{U}_{\text{sep}}$ is the separation between units (see Fig. S2).
A.2.4 Static and Dynamic Response
@@
Using the parameters from the fitting process, we can predict stable state configurations for the DPF with the resulting model. Given that the non-linear elements of the gripper are independently stable, the expected number of states for a given dome phalanx finger topology follows $2^{n}$, where $n$ is the number of bistable segments on the finger. The model can determine the behavior of a given design in orders of magnitude less time than the FE numerical simulations, which makes it feasible for iterating through potential configurations and optimizing for the best design according to a given task (e.g., position and grasping force).
Static response: The static response of the finger is predicted by minimizing the total energy of the geometry. We performed this process using Matlab’s fmincon sqp algorithm [53]. Our algorithm can predict different final shapes depending on the initial guess geometry. Results for the model are compared with FE simulations, where various points of the limiting layer are compared (see Fig. S5a). Results show that our model can predict the final shape of the DPF for different geometric parameters with an error below 10% compared with FE simulations.
Dynamic response: The model can be further expanded by including point masses on each of the segments to represent the total mass of the system and dampers to capture the visco-elastic response of the material. By incorporating these elements and using symbolic differentiation, a dynamic model can be constructed to capture the time response of the DPF. The equations of motion for the system can be written as
$[M]\ddot{q}+[D]\dot{q}+\nabla\left(E_{L}+E_{NL}+E_{T}\right)=F_{ext}(t)$ (S.17)
Where $[M]$ is the mass matrix of the system, $[D]$ is the damping due to viscous effects, and $F_{ext}$ is the external actuation pressure. By substituting $\dot{x}_{1}=x_{2}$ and $\dot{x}_{2}=\ddot{q}$, we can rewrite the system as
$\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}x_{2}\\ -[M]^{-1}\left(\nabla\left(E_{L}+E_{NL}+E_{T}\right)-[D]q+F_{ext}(t)\right)\end{bmatrix}$ (S.18)
The system is solved using the Runge-Kutta time integration method for loading and unloading external force conditions. Time response for different DPF geometries can be observed in Fig. S5c.
@@
Figure S5: Static and dynamic response of the DPF. a) Stable states are found by minimizing the strain energy. A comparison with FE simulations for five domes of the same height $\rightarrow$ H = 4 mm and H = 5 mm. b) Dynamic response of lattice model with visco-elastic dampers c) Five segment DPF dynamic response for different dome H.
A.3 Geometry Optimization
A.3.1 Inverse problem $\rightarrow$ Position Optimization
@@
The final results for the position optimization can be observed in Fig. S6, where the five target positions and the final shape predicted by the model are shown. The final geometry and objective function of the optimization algorithm are summarized in Table. S1, where we can observe that for targets 1,2 and 5, the desired position is attained with almost perfect precision.
@@
Table S1: Inverse design problem results for five different targeted positions.
@@
UC 16.0 16.0 16.0 16.0 16.00
Obj 0.2 0.2 2.9 2.7 0.4
A.3.2 Inverse problem $\rightarrow$ Position + Stiffness Optimization
We further expand our model to include stiffness optimization in our objective function. By incorporating finger stiffness, grasping force is increased by changing the geometric parameters. Results for the optimization algorithm, where the same number of segments as in Table S1 is maintained, are shown in Table S2.
Table S2: Inverse design problem results for five different targeted positions with maximum stiffness.

### v2

@@ A.3.3 Parameter Tuning
To determine each of our lumped model’s parameters, the DPF’s fundamental unit cell is isolated into two different parts (see Figure S5a). The dome unit is fitted as a nonlinear spring following Equation 1 and the strain limiting layer is represented by a liner spring to capture stretching and two rotational springs to capture the bending rigidity. Each of the constants for these springs needs to be tuned for different geometrical cases to produce an accurate model for inverse co-design. Given Equation 1, the parameters to be determined for the nonlinear spring are $k_{b}$, $\alpha$, and $d$. These parameters are functions of the dome unit’s mechanical behavior (stiffness and energy barrier) as well as the interaction between the dome and the wall of the air chamber of the subsequent unit cell. The procedure to find each constant as a function of the geometrical parameters of the dome unit cell is as follows. First, different finite element simulations of the unit cell are utilized to determine the best-fitted values for $k_{b}$, $\alpha$, and $d$ (Figure S5b) for different geometrical cases ($H$ and $t$, same as shown in Figure S1). Different parameters, such as elastic modulus, are examined to establish their influence on the strain energy and tip displacement (Figure S5b) and establish a better understanding of the governing parameters to be considered to represent the nonlinear spring accurately. Equation 1 is fitted to the FE data using the nonlinear least square method to obtain the overall behavior of these parameters in terms of $H$ and $t$. Results for these fitted parameters can be observed in Figure S5c, and the energy-displacement curve with the fitted parameters can be observed in Figure S6a.
@@
Figure S5: a) DPF fundamental unit cell isolated into different lumped parameter spring elements. (nonlinear + linear + rotational springs) b) nonlinear spring parameter behavior over different dome heights $H$ and dome thickness $t$.
@@
Once the values are obtained, we use Recursive Feature Elimination (RFE) [50] combined ridge regression to determine an expression for $k_{b}$, $\alpha$, and $d$ as functions of the dome height ($H$), dome thickness ($t$), and dome curvature ($R$). By using RFE, we can automatically determine the relevance of each feature in our model. Given the main task of inverse design of multistable actuators, a general and scalable model is needed to cover different design cases. We establish the model by considering three different non-dimensional relations:
@@
where superscripts $n$ and $m$ represent the maximum degree of the polynomial use for the regression and subscripts $i$ and $j$ represent each of the non-dimensional numbers. For the parameter $\alpha$ we can fully represent this regression as:
$\begin{split}&\begin{bmatrix}\alpha(H_{1},t_{1},R_{1})\\ \alpha_{2}(H_{2},t_{2},R_{2})\\ \vdots\\ \alpha_{k}(H_{k},t_{k},R_{k})\end{bmatrix}=\\ &\begin{bmatrix}\pi_{i}(H_{1},t_{1},R_{1})&\ldots&\pi^{n}_{i}(H_{1},t_{1},R_{1})&\pi_{i}\pi_{j}(H_{1},t_{1},R_{1})&\ldots&\pi^{n}_{i}\pi^{m}_{j}(H_{1},t_{1},R_{1})\\ \pi_{i}(H_{2},t_{2},R_{2})&\ldots&\pi^{n}_{i}(H_{2},t_{2},R_{2})&\pi_{i}\pi_{j}(H_{2},t_{2},R_{2})&\ldots&\pi^{n}_{i}\pi^{m}_{j}(H_{2},t_{2},R_{2})\\ \vdots\\ \pi_{i}(H_{k},t_{k},R_{k})&\ldots&\pi^{n}_{i}(H_{k},t_{k},R_{k})&\pi_{i}\pi_{j}(H_{k},t_{k},R_{k})&\ldots&\pi^{n}_{i}\pi^{m}_{j}(H_{k},t_{k},R_{k})\end{bmatrix}\begin{bmatrix}\xi_{1}\\ \xi_{2}\\ \vdots\\ \xi_{k}\end{bmatrix}\end{split}$ (S.15)
Figure S6: Parameter-fitting results for $\alpha$, $k_{b}$, $d$ as a function of geometric parameters. a) RFE with ridge regression cross-validation vs the number of features (training data) for $k_{b}$ (i), $\alpha$ (ii), and $d$ (iii); b) Prediction vs observation plot on test data for the reduced model for $k_{b}$ (i), $\alpha$ (ii), and $d$ (iii); c) Model fit for energy as a function of geometric parameters for monostable (i), metastable (ii) and bistable (iii) cases.
Note that we assume that the library of terms used to obtain $\mathbf{C}(\pi_{1},\pi_{2},\pi_{3})$ has a sufficiently rich column space that the behavior will be represented by Equation S.13, and it can be written as a linear combination of the weights $\xi$. The data is split into 70% train and 30% test data sets to calculate the accuracy of our model. Observation vs prediction plot and estimation error for the three different variables can be observed in Figure S6b (ii), where a good fit between the proposed model and the test data is observed. Moreover, we use RFE with cross-validation (cv) on the test data to eliminate features and retest the model to determine its accuracy. $R^{2}$ for the cross-validation data as a function of the number of features can be observed in Figure S6b(i). This gives a reduced model with similar prediction capabilities and the most relevant geometric relations of the problem, and yields an accurate representation of the behavior (see Figure S6b (iii)).
@@
$\begin{split}k_{b}=f\left(E,\pi_{1},\pi_{2},\pi_{3}\right)=&\frac{E}{R^{2}}\Bigl{(}-\frac{1.97Rt^{5}}{H^{3}}+\frac{7.4H^{2}t^{3}}{R^{2}}+\frac{3.5Rt^{4}}{H^{2}}+\frac{0.37H^{2}t^{2}}{R}+\frac{42.2t^{5}}{H^{2}}\\ &-\frac{35.8Ht^{4}}{R^{2}}+\frac{71.8t^{5}}{HR}-\frac{3.4Rt^{3}}{H}-\frac{67.7t^{4}}{H}+4.2Rt^{2}+11.1t^{3}\Bigr{)}\end{split}$ (S.16)
$\begin{split}\alpha=f\left(\pi_{1},\pi_{2},\pi_{3}\right)=&\frac{0.29H^{3}}{R^{3}}-\frac{0.44t^{3}}{H^{3}}-\frac{2.2H^{2}t}{R^{3}}-\frac{0.73H^{2}}{R^{2}}-\frac{6.9t^{3}}{H^{2}R}+\frac{1.17t^{2}}{H^{2}}-\frac{3.0t^{3}}{HR^{2}}+\\ &\frac{7.05Ht}{R^{2}}+\frac{13.38t^{2}}{HR}+\frac{0.44H}{R}+\frac{13.67t^{3}}{R^{3}}-\frac{15.0t^{2}}{R^{2}}-\frac{4.7t}{R}\end{split}$ (S.17)
$d=f\left(\pi_{1},\pi_{2},\pi_{3}\right)=2.14H+0.25t-H-\text{U}_{\text{sep}}-t_{ch}/2-t/2$ (S.18)
@@
where $H$, $t$ and $R$ are dome unit parameters, $E$ is the elastic modulus of the material,$\text{U}_{\text{sep}}$ and $t_{ch}$ is the separation between units (see Figure S2).
A.4 Static and Dynamic Response
@@
Using the parameters from the fitting process, we can predict stable state configurations for the DPF with the resulting model. Given that the nonlinear elements of the gripper are independently stable, the expected number of states for a given dome phalanx finger topology follows $2^{n}$, where $n$ is the number of bistable segments on the finger. The model can determine the behavior of a given design in orders of magnitude less time than the FE numerical simulations, which makes it feasible for iterating through potential configurations and optimizing for the best design according to a given task (e.g., position and grasping force).
@@
Figure S7: Static and dynamic response of the DPF. a) Stable states are found by minimizing the strain energy. A comparison with FE simulations for five domes of the same height. b) Dynamic response of lattice model with viscoelastic dampers c) Schematic representation of applied load and final DPF position. d) Five-segment DPF dynamic response for different dome H.
Static response: The static response of the finger is predicted by minimizing the total energy of the geometry. We performed this process using Matlab’s fmincon sqp algorithm [57]. Our algorithm can predict different final shapes depending on the initial guess geometry. Results for the model are compared with FE simulations, where various points of the limiting layer are compared (see Figure S7a). Results show that our model can predict the final shape of the DPF for different geometric parameters with an error below 3% compared with FE simulations (see Table S1 for an extended comparison).
Dynamic response: The model can be further expanded by including point masses on each of the segments to represent the total mass of the system and dampers to capture the viscoelastic response of the material. By incorporating these elements and using symbolic differentiation, a dynamic model can be constructed to capture the time response of the DPF. The equations of motion for the system can be written as
$[M]\ddot{x}_{i}+F_{d}^{int}(x_{ij},\dot{x}_{i},\dot{x}_{j})+F_{d}^{iso}(\dot{x}_{i},\dot{x}_{j})+\nabla\left(E_{L}(x_{ij})+E_{NL}(x_{ij})+E_{T}(x_{ij})\right)=F_{ext}(x_{i},t)$ (S.19)
where $[M]$ is the mass matrix of the system, $F_{d}^{int}(x_{ij},\dot{x}_{i},\dot{x}_{j})$ is the internal damping, which can be modeled as a dashpot in parallel to each spring (see Figure 3a) and $F_{d}^{iso}(\dot{x}_{i})$ is the isotropic drag, which acts as an external force on each node proportional to velocity ($\dot{x}_{i}$[55]. $F_{ext}(x_{i,j},t)$ is the external actuation pressure that can be calculated as $F_{ext}(x_{ij},t)=A(x_{ij},)P(t)$. Where $A(x_{ij},)$ is the area of the face where the pressure is applied. By substituting $\dot{x}_{1}=x_{2}$ and $\dot{x}_{2}=\ddot{x}_{i}$, we can rewrite the system as
$\begin{bmatrix}\dot{x}_{1}\\ \dot{x}_{2}\end{bmatrix}=\begin{bmatrix}x_{2}\\ -[M]^{-1}\left(F_{d}^{int}(x_{ij},\dot{x}_{i},\dot{x}_{j})+F_{d}^{iso}(\dot{x}_{i},\dot{x}_{j})+\nabla\left(E_{L}(x_{ij})+E_{NL}(x_{ij})+E_{T}(x_{ij})\right)-F_{ext}(x_{ij},t)\right)\end{bmatrix}$ (S.20)
The system is solved using the Runge-Kutta time integration method for loading and unloading external force conditions. We explore the dynamic behavior of a five-segment DPF by applying the internal pressure as shown in Figure S7c, where we obtain the final stable state as shown and the dynamic response shown in Figure S7d. As expected, the model can capture the dynamic response of a bistable, metastable, and monostable DPF, where snap-through and snap-back phenomena are observed.
Furthermore, we analyze the temporal response of the metastable units to assess how geometrical parameters and material damping influence the resetting time. Our results show that adjusting the dome unit’s thickness modifies the resetting time (see Figure S9b). However, an even more pronounced effect is observed when varying the loading duration (see Figure S8c). This suggests that after manufacturing the DPF, additional tunability can be achieved by controlling the applied pressure (see Movie 7).
Figure S8: Lattice model dynamic response for dome unit cell. a) Internal pressure over time $P(t)$. Loading time $\tau_{1}$ and applied load time $\tau_{2}$. b) Effect of internal damping on dynamic response. c) Effect of applied load time ($\tau_{2}$) on metastable unit dome reseating time.
The resetting time response is validated against experimental data to assess its predictive accuracy. As shown in Figure S9c, the model effectively captures the time scales of a fully metastable DPF. However, due to the inherent randomness in the resetting process when all units are metastable, the tip displacement is not precisely predicted. Nevertheless, this tool enables the design of both the static and dynamic responses by programming the resetting time scales of the metastable domes (see Movie 2 for experimental comparison).
Figure S9: Metastable dome unit analysis. a) Effect of dome height in resetting time of the units (t = 0.8 mm). b) Effect of dome thickness in resetting time of the units (H = 3.0 mm). c) Comparison between experimentally recorded data shown in Figure 2c and our proposed model.
A.5 Model Validation and Simulation time
The lattice model is evaluated under different geometric configurations as shown in Table S1.
N Segments $R_{b}$ $H/R_{b}$ t E $t_{\text{lim}}$ Error %
1 5 5 0.7 0.7 5 1.4 1.37
2 5 5 0.8 0.8 5 1.6 0.62
3 5 5 1 1 26 2 0.51
4 5 8 0.7 0.75 5 1.5 0.51
5 5 8 0.8 1 26 1.6 2.73
6 6 5 0.7 0.7 5 1.4 1.42
7 6 5 0.8 0.8 5 1.6 0.6
8 6 5 1 1 26 2 0.4
9 6 8 0.7 0.75 5 1.5 0.44
10 6 8 0.8 1 26 1.6 2.47
Table S1: List of validation cases for spring model. Error % reference to the difference between model and FE analysis.
Different numbers of segments, elastic modulus, and limiting layer thickness are included to demonstrate the versatility of our approach. The error between the model and the simulation is calculated by comparing the difference between the final position of the limiting layer after all units are activated (see Fig S10).
Figure S10: Model comparison with FE method. (1)-(10) Case numbers parameters and error can be observed in Table S1.
Moreover, our model computational cost is evaluated by comparing the total simulation time simulation time with FE simulations. Results show that the model can yield the final shape of the DPF in less than 1 s. More importantly, the model does not exhibit a coupling between the number of segments and computational time. (see Table S2).
N segments FE Time [s] FE Time [h] Model Time [s]
2 3094 0.86 0.66
3 5960 1.66 0.58
4 12072 3.35 0.64
5 15859 4.41 0.76
6 21455 5.96 0.67
Table S2: Simulation time comparison between FE analysis and spring model.
A.6 Geometry Optimization
A.6.1 Inverse problem $\rightarrow$ Position Optimization
@@
The final results for the position optimization can be observed in Figure S11, where the five target positions and the final shape predicted by the model are shown. The final geometry and objective function of the optimization algorithm are summarized in Table S3 and S4, where we can observe that for targets 1,2 and 5, the desired position is attained with almost perfect precision.
@@
Table S3: Inverse design problem results for five different targeted positions (NinjaFlex 85A).
@@
t 0.6 0.62 0.68 0.62 0.75
$t_{\text{lim}}$ 1.5 1.3 1.2 1.4 1.25
UC 15.0 15.0 15.0 15.0 15.0
$K_{\text{dis}}$ 0.009 0.025 0.01 0.018 0.007
Obj 0.2 0.2 2.9 0.8 0.4
Table S4: Inverse design problem results for five different targeted positions (Cheetah 95A).
Target 1 Target 2 Target 3 Target 4 Target 5
[46 32] [25 20] [25 40] [25 30] [35 45]
H1 4.87 4.52 4.89 4.96 4.9
H2 4.8 3.93 4.72 4.52 4.87
H3 4.67 3.63 4.52 4.08 4.72
H4 3.98 4.33 3.99 4.35
H5 4.02 4.3
H6
Unit S 2.31 1.10 1.17 1.08 1.52
Unit L 9.83 7.28 6.95 6.93 8.8
t 0.79 0.83 0.65 0.76 0.63
$t_{\text{lim}}$ 1.63 1.59 1.19 1.07 1.72
UC 15.0 15.0 15.0 15.0 15.0
$K_{\text{dis}}$ 0.093 0.341 0.079 0.112 0.063
Obj 0.775 0.056 0.55 0.22 0.32
A.6.2 Inverse problem $\rightarrow$ Position + Stiffness Optimization
We further expand our model to include stiffness optimization in our objective function. By incorporating finger stiffness, grasping force is increased by changing the geometric parameters. Results for the optimization algorithm, where the same number of segments as in Table S3 and S4 is maintained, are shown in Table S5 and S6.
Table S5: Inverse design problem results for five different targeted positions with maximum stiffness (NinjaFlex 85A).